A buffer is prepared containing 3.00 molar acetic acid and 1.00 molar sodium acetate. What is its pH? The pKa of aetic acid = 4.75
Answer #1

It is given that, Concentration of acetic acid = [CH3COOH) = 3.00 M Concentration of sodium acetate = [CH3COONa] = 1.00 M pkapH = pka + log[CH3COONa] / [CH3COOH] pH = 4.75 + log[1.00 M] / [3.00 M) = 4.75 -0.477 pH = 4.273